OpenStudy (anonymous):

let \( \Large \vec{F(r)}=\frac{c\vec{r}}{r^3} \) , \(\large \vec{r} = x\vec{i}+y\vec{j}+z\vec{k} \), Prove \( \Large \iint_{S} \vec{F}.d\vec{S} \), which \( S \) is a sphere, does not depend on the radius of the sphere.

OpenStudy (anonymous):

it's not a hard problem, but i'm looking for some innovative and easy solutions ;)

OpenStudy (irishboy123):

divergence theorem nails it

OpenStudy (anonymous):

\(\Large \iint_{S} \vec{F}.d\vec{S} = \iiint_{E} div\vec{F}dV\), therefore, \(\Large \iint_{S} \vec{F}.d\vec{S}=\frac{ 3c}{a^3} \iiint_{E} dV =4 \pi c \)

OpenStudy (anonymous):

@IrishBoy123 , correct ;) how about using Stokes' thm or other ideas?